OpenStudy (anonymous):

Ms. Green tells you that a right triangle has a hypotenuse of 13 and a leg of 5. What is the length of the other leg?

OpenStudy (mathstudent55):

\(a^2 + b^2 = c^2\) a and b are the legs; c is the hypotenuse.

OpenStudy (anonymous):

I got 194 how do I simplify that to the answer then square?

OpenStudy (anonymous):

this triangle is a special type of triangle called a pythagorean triple... there are not many but easy to memorize... which is a 5, 12, 13 triangle... but if you want to solve you have a and b as legs and c as hypotenuse. 5^2+b^2=13^2

OpenStudy (anonymous):

and then just solve for b.

OpenStudy (anonymous):

8?

OpenStudy (mathstudent55):

Since the hypotenuse is 13, the square of 13 is 169. The leg is 5. The square of 5 is 25. The squares of the legs are added together. If you have the hypotenuse and one leg, subtract the square of the leg from the square of the hypotenuse. 169 - 25 = 144 144 is the square of the other leg. 144 is the square of 12.

OpenStudy (anonymous):

no... you have 25+b^2=169... and then subtract 25 from both sides... and you get b^2=144... then take the square root of 144, which is 12... which like i said is a special kind of triangle called a pythagorean triple... 5, 12, 13 triangle.